OpenStudy (anonymous):

Jay finds a texting plan that costs a flat fee of $15 per month for 100 texts, plus $.25 per text after that. Which cost function represents this scenario if x equals the number of texts exceeding 100 that Jay sends in a month? C(x) = 100 + .25x C(x) = .25+ 100x C(x) = .25 + 15x C(x) = 15+ .25x

OpenStudy (anonymous):

Im stuck on these, its been so long since ive done them, I cant remember how to solve them, but if I had to take a guess I would say for the second one, A but I have no idea for the first one, still figuring it out...

OpenStudy (anonymous):

which question do you want to do 1st?

OpenStudy (anonymous):

we can do the first one, but sadly your going to have to explain it to me... I haven't done these kinds of problems in a while so i'm not to familiar with it :( sorry

OpenStudy (anonymous):

ok. c(x) if function notation for the equation or function that will compute the monthly cost as described in the question. ok thus far?

OpenStudy (anonymous):

plus, im probably going to have a lot of questions... this pre test for my final has material that I learned a long time ago, and its not even what I learned in this semester... so sadly im gonna have to go through it all over again... I dont know if ur up to help me that much but I would truly appreciate any kind of help! :)

OpenStudy (anonymous):

okay, I got that so far...

OpenStudy (anonymous):

ok. so what is he going to pay for sure each month no matter how many texts?

OpenStudy (anonymous):

$15 plus $.25

OpenStudy (anonymous):

no. $15. if he texts 37 or 21 or 0 times, he is still going to pay $15. only when he goes over 100 texts will he pay more. so he pays $15 whether he texts or not. ok?

OpenStudy (anonymous):

ohh okay!

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

so C(x) must equal 15 plus whatever additional charges he incurs. so the additional would be added to 15. what does the question state about what x represents?

OpenStudy (anonymous):

x represents the number of texts exceeding 100 that Jay sends in a month

OpenStudy (anonymous):

so if he texts 110 times a month, then x=10, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what does 10 extra texts cost him?

OpenStudy (anonymous):

would it be 2.5, because .25(10) = 2.5?

OpenStudy (anonymous):

correct. you multiplied the rate--.25--times the number of texts over 100 (which is represented by x as stated in the question. so which answer do you think is the correct one?

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

correct. so why do you think this is so ahrd?

OpenStudy (anonymous):

It not hard, I just didn't remember how to solve it... lol thanks for your help! but I do have some more questions... can you stick around?

OpenStudy (anonymous):

but I have to go to the restroom really quick ill brb, so please stick around :)

OpenStudy (anonymous):

i can at least for the graph question posted

OpenStudy (anonymous):

okay, im back and cool, thanks :)

OpenStudy (anonymous):

so, u want to get started on the graph one?

OpenStudy (anonymous):

yes. to solve ineqs, we use the same method as for equations--isolate x by algebraic manipulation--following the usual rules WITH ONE MAJOR EXCEPTION: WHEN YOU MULTIPLY OR DIVIDE BY A NEGATIVE NUMBER, YOU MUST SHIFT THE DIRECTION OF THE INEQ SIGN. SO WHAT do you come up with?

OpenStudy (anonymous):

okay, so I have −5 < − 4x + 3 ≤ 23 can u remind what to do from here?

OpenStudy (anonymous):

you have to have x by itself in the middle. so start by subtracting 3 FROM EACH MEMBER OF THE INEQ. what do you then have?

OpenStudy (anonymous):

okay from there you would get: −5 < − 4x + 3 ≤ 23 -3 -3 -3 -8< -7x ≤ 20 you would get this right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay, can you explain...

OpenStudy (anonymous):

everything is right except when you subtracted 3 from -4x+3 you don't get 7x, you get -4x--you subtract the 3 from the 3 and not the -4x

OpenStudy (anonymous):

ohhh, okay, so its -8 < -4x ≤ 20

OpenStudy (anonymous):

right. Now you have to still get x by itself

OpenStudy (anonymous):

combine like terms?

OpenStudy (anonymous):

no. x is "modified" by a "-4" so you need to divide everything by -4

OpenStudy (anonymous):

ohh right! okay okay... -8/-4 < -4x/-4 ≤ 20/-4 = 2 < 1x ≤ -5? not sure about the middle part...

OpenStudy (anonymous):

you forgot the MAJOR EXCEPTION TO THE RULES: you divided by a negative number so you have to change the dirction of the ineq signs. ( and 1x is written simply as "x")

OpenStudy (anonymous):

ohh RIGHT!!! ugh sorry the answer would be 2 > 1x -> -5

OpenStudy (anonymous):

right. so x is less than 2 but greater than or equal to -5. Now look at the first choice for the graphic answer. The shaded portion of the number line represents all the numbers on the shade portion. The open circle on 2 means that 2 is not include and the colored circle on the -5 means that -5 is included in the solution. so do you think that this one is the correct graph?

OpenStudy (anonymous):

yes, I think A is the correct one!

OpenStudy (anonymous):

you are correct

OpenStudy (anonymous):

yay!! :) wow thank you so much for going over this with me!! I do have some more... do you think you can stick around?

OpenStudy (anonymous):

I have these three, two of which I have answered, but im stuck on the middle one... 6. Maria brings d dollars to an amusement park. The park charges an admission fee of $20. Each game at the park costs $1.50. Which inequality shows the number of games, g, which Maria can play? (4 points) 20 + 1.5g ≥ d 20 + 1.5g < d 20 + 1.5g > d 20 + 1.5g ≤ d<<<My Answer 7. Which equation represents the line passing through the points (2, −2) and (−5, 12)? 2x − y = 2 2x + y = 2 x + 2y = 2 x − 2y = 2 8. A company makes jewelry. It costs $6,400 to make 10 pieces of jewelry and $11,200 to make 20 pieces of jewelry. Which equation models the cost, C(x), as a linear function of the number of pieces of jewelry, x? C(x) = 480x + 1600<<< My Answer C(x) = 1600x + 480 C(x) = 480x − 1600 C(x) = 1600x − 480

OpenStudy (anonymous):

and I apologize if im overwhelming you, I just really want to be ready to take my final, so im really stressing over it...

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

Here are some links to some more of my questions, now take note that the same ones we just went over are on here as well, so pay no mind to them, just the ones you haven't seen... http://answers.yahoo.com/question/index;_ylt=Ahp4jdZwsHodeE5Ns9k1hhnsy6IX;_ylv=3?qid=20120504084358AAWAaLt http://answers.yahoo.com/question/index;_ylt=AotJ07oO.0iNL.Ty7i4mTSTsy6IX;_ylv=3?qid=20120504090545AACpTly

OpenStudy (anonymous):

for 7, plug in the 2 points into each eq and see if it is true. The links say that I can't access.

OpenStudy (anonymous):

try copying the whole thing, mite help

OpenStudy (anonymous):

come to my new question, it will be easier